OpenStudy (anonymous):

the circular blade on a saw has a diameter of 7.5 inches and rotates at 2400 revolutions per minute. A) find the angular speed in radians per sec B) Find the linear speed of the saw teeth ( in feet per sec) as the contact the wood being cut. Formulas--> Linear speed= arc length/time= s/t Angular speed= central angle/time

OpenStudy (anonymous):

I'm not sure if I'm right but. Angular speed is 4800 Rad/Second Linear velocity is 7.5 *2*Pi*2400=1.13097... × 10^5 RAD Not sure

OpenStudy (anonymous):

\[4800\pi RAD/Second\]

OpenStudy (anonymous):

the answer in the book is 80pi rad per sec for a and 25pi ft per sec for b but i dont know how to get to that...

OpenStudy (anonymous):

One minute I will get this done.

OpenStudy (anonymous):

(2400*2*PI)/60(seconds) = 80 PI RAD/Sec

OpenStudy (anonymous):

this formula might help --> arc length s= radius r * central angle(in radians)

OpenStudy (anonymous):

it gives me 25 only if I do ((80*pi*7.5)/(2*pi)) and all that convert to feet it will be exatly 25

OpenStudy (anonymous):

but as in our case if the angular velocity is 80pi Rad/s should be V = 80pi *7.5 which give us 157. Sorry

OpenStudy (anonymous):

its okay but thanks for helping me! :)